OpenStudy (anonymous):

Write an equation of a line that passes through the point (0, 4) and whose slope is 4. ?????

OpenStudy (aroub):

y=mx+b , m=4 , (0,4) 4=4(0)+b 4=b y=4x+4

OpenStudy (anonymous):

bingo!

OpenStudy (aroub):

:D